1. the differential equation(1+2"x" )"dy\/dx" +"xy" =0 , determine the particular solution when "y" (0)=2, express y interms of "x" .
  2. show that "dy\/dx="e^(y-x) is seperable and find the general solution expessing y in terms of x
  3. use the U substitution "y=tx" to solve "(x^6-2x^5+2x^4-y^3+4x^2y)dx+(xy^2-4x^3)dy=0"
1
Expert's answer
2021-10-20T17:58:55-0400

2.

"dy\/e^y=dx\/e^x"

"-e^{-y}=-e^{-x}+c"

"y=-ln(e^{-x}-c)"


1.

"dy\/y=-xdx\/(1+x^2)"

"lny=-ln(1+x^2)\/2+lnc"

"y=c/\sqrt{1+x^2}"


"y(0)=c=2"


"y(x)=2/\sqrt{1+x^2}"


3.

"y=tx"

"(x^3-2x^2+2x-t^3+4t)dx+(t^2-4)(xdt+tdx)=0"

"(x^3-2x^2+2x)dx+(t^2-4)xdt=0"

"\int(x^2-2x+2)dx=-\int(t^2-4)dt"


"t^3\/3-4t=-x^3\/3+x^2-2x+c"


"y^3\/(3x^3)-4y\/x=-x^3\/3+x^2-2x+c"